OpenStudy (anonymous):

How would I simplify these questions with negative integers? 3a / 4a^-4 ; 2x^-2 / 3x^4 ; ( (2^0)^2/ 2^3 x 2^3 )^-3 ; 8x^-2 / x^-4

geerky42 (geerky42):

HINT: \(\large a^{-b} = \dfrac{1}{a^b}\)

geerky42 (geerky42):

\[\large \dfrac{1}{a^{-b}} = a^b\]

geerky42 (geerky42):

For example, \(\large 3^{-4} = \dfrac{1}{3^4}\)

geerky42 (geerky42):

Does this help?

OpenStudy (anonymous):

^ Thank you. It does :)